OpenStudy (anonymous):

A map has a scale of 1 cm : 11.5 km. Two cities are 5.5 cm apart on the map. To the nearest tenth of a kilometer, what is the actual distance corresponding to the map distance?

OpenStudy (texaschic101):

proportion time... 1/11.5 = 5.5/x (1 cm to 11.5 km = 5.5 cm to x km) cross multiply (1)(x) = (11.5)(5.5) x = 63.25 km